OpenStudy (anonymous):

A study of social mobility in America examined the social class attained by the sons of lower class fathers. Social classes were numbered from 1 to 5 with 1 representing the lower class and 5 the higher class. Consider the random variable X to the class of a randomly chosen son. The study found the following distribution: Son's Class 1 2 3 4 5 Probability 0.48 0.38 0.08 0.05 0.01 a) What percent of the sons reached the highest class? 0.01 b) Check that this distribution meets the requirements of a discrete probability distribution. c) What is P(X < 2)? 0.48 d) What is P(X = 2)?0.38 or 38% e) Write the event: a son of a lower-class father attains one of the highest two social classes in terms of X. (6 points) P(X>4)=(PX=4)=P(X-5) 0.05=0.01=0.06

ganeshie8 (ganeshie8):

for part a) it should be 1%

ganeshie8 (ganeshie8):

0.01 means 1% okay ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

wat about part b ?

ganeshie8 (ganeshie8):

u have done a similar thing in ur previous problem

OpenStudy (anonymous):

0.48+0.38+0.08+0.05+0.01=1

OpenStudy (anonymous):

Its satisfies the 2 requirements

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

part c is right !

ganeshie8 (ganeshie8):

part d is also right

ganeshie8 (ganeshie8):

for part e) : P(X >= 4) = P(X = 4) + P(X = 5) = 0.05 + 0.01 = 0.06